# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v1

@@ Acknowledgements
The author thanks T.Drivas for pointing out the ongoing work [DG26+] and for interesting discussions. The author thanks A.Mazzucato and A.Novikov for stimulating discussions and valuable suggestions. A. Pan was partially supported by the US National Science Foundation grants DMS-1909103 and DMS-2206453 (PI Mazzucato).
@@ 2 Geometry of $\mathsf{SDiff}(D)$ with Weighted Metric
In this section, we study the Riemannian structure of $\mathsf{SDiff}(D)$ which naturally fits the inhomogeneous incompressible Euler equation. Our presentation in this section mainly follows and generalizes the discussions in the unpublished note [DrWS25], which generalizes constrained geodesic, second fundamental form and sectional curvature of level set submanifold from finite to infinite dimension (The $\mathsf{SDiff}(D)$ setting).
We start with the well-known Jacobi formula which will be repeatedly used in this section.
@@ 2.1 The Riemannian Structure
Let $D$ be either $\mathbb{R}^{d}$ or $\mathbb{T}^{d}$, which is the physical domain of our fluids, with natural decaying/periodic boundary conditions. Throughout this paper, we denote by $\mathsf{Diff}(D)$($\mathsf{SDiff}(D)$) the Lie group of (volume-preserving) diffeomorphisms in $D$. We denote by $\mathfrak{X}(D)$ ($\mathfrak{X}_{\sigma}$) the space of (divergence-free) smooth vector fields on $D$. We remark that the formulation below works in the setting of $H^{s+1}$ Sobolev regularity, with $s>d/2$. Here, we only consider the smooth case for simplicity.
@@
here $\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A$. Clearly $\rho$ is a rearrangement of $\rho_{0}$, so the assumption (A) is preserved under action of $\mathsf{SDiff}(D)$ on $\mathsf{Dens}(D)$.
@@
We now extend the domain of the above normal generator $\mathcal{N}$ to distributions. Let $\mathcal{M}$ be the set of Borel measures with finite variation, for any $\mu\in\mathcal{M}$ we define the following co-normal functional
@@
which satisfies the annihilating property $T_{X}\mathsf{SDiff}(D)\subset\mathsf{Ker}(\mathcal{N}_{\mu}^{X})$.
@@
    where $\Delta_{\rho}=\nabla\cdot(\rho^{-1}\nabla)$ and $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$.
@@ 2.3 Second Fundamental Form and Curvature Tensor
$\int_{D}[\mathbf{Q}_{\rho}(u\cdot\nabla u)\mathbf{Q}_{\rho}(v\cdot\nabla v)-\lvert\mathbf{Q}_{\rho}(u\cdot\nabla v)\rvert^{2}]\rho(z)dz$
@@
$\ddot{X}(t,a)=(\partial_{t}u+u\cdot\nabla u)(X_{t}(a))=0$
which implies that streamlines are all straight lines.
@@ 2.4 A Gibbs-Appell Formulation
Different from the celebrated Lagrange’s least action principle/Hamilton’s stationary action principle in analytical mechanics, there is a much lesser popular principle/formalism in classic mechanics, known as Gibbs-Appell formulation, which connects to the so-called Gauss’s principle of least curvature and Hertz’s principle of least constraint (see [[Lew96]]). The natural playground for this formulation is a mechanical system with ideal constraints, which is compatible with the nature of incompressible fluids. We refer to [[TG23]] for the study of the Navier-Stokes equation from this point of view. Instead of minimizing Lagrangian action under the constraint of fixed endpoints, the Gibbs-Appell principle minimizes the so-called Appellian functional, which is a quadratic function in acceleration, at each instant.
@@
Then minimizer $X_{*}$ of $\mathscr{A}$ at each instant $t\in[0,T]$, subject to the constraint $X\in\mathsf{SDiff}(D)$, is the solution of the following Newton’s equation:
@@ 3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
We now state the following:
@@
    Given initial data $(u_{0},\xi_{0})$ satisfying the compatibility condition, the solution of (30) is given by the following Lagrangian-Eulerian system:
@@ 3.2 Vorticity Formulation
We formally unify the above relations in dimension $2$ in the following commutative diagram.
@@
$\mathbf{P}_{\rho}(\rho^{-1}\eta)=\nabla^{\perp}\Delta_{\rho}^{-1}\nabla\times\eta.$
@@
Assume $(u,\rho)$ is a smooth solution of IIE, then for any regular value $\alpha>0$ of $\rho_{0}$, where $E_{\alpha}:=\{\rho_{0}=\alpha\}$ is a smooth loop. Let $D_{\alpha}:=\{\rho_{0}<\alpha\}$ is a smooth bounded domain enclosed by $E_{\alpha}$, we have
@@ 3.3 Local Existence Via Vorticity Formulation
Assume $\omega\in H_{0}^{2}$, $\rho\in H^{3}$ satisfying the assumption (A), then any mean-zero strong solution $\psi$ of $\mathbf{L}_{\rho}\psi=\omega$ also satisfies $\psi\in H_{0}^{4}$.
@@
Step 1.($H^{1}$ Estimate) Multiplying both sides of the equation by $\psi$ and integrating by part, we have:
@@
Now, since $\rho\in H^{3}$ and the function $f(z)=z^{-1}$ is smooth on the range of $\rho$, we have $\rho^{-1}=f(\rho)\in H^{3}$. Hence, we conclude:
$\lVert\psi\rVert_{H^{3}}\lesssim\lVert\omega\rVert_{H^{1}}\lVert\rho^{-1}\rVert_{H^{2}}(1+\lVert\rho\rVert_{H^{3}}+\lVert\rho\rVert_{H^{3}}^{2})\lesssim\lVert\omega\rVert_{H^{1}}\sum_{j=1}^{3}\lVert\rho\rVert_{H^{3}}^{j}$
@@
where $P$ is some polynomial of degree $5$ with positive coefficients, hence it is monotonically increasing on $(0,\infty)$. ∎
We now prove the local existence of IIE using our vorticity formulation.
@@
Let $(\omega_{0},\rho_{0})\in H^{2}\times H^{3}$ and let $\rho$ satisfies Assumption (A), then (47) admits a unique solution local in time.
@@
Now we apply $J^{2}:=1-\Delta$ to the vorticity equation, pairing with $J^{2}\omega^{n}$ and integrating by part we have:
$\frac{d}{dt}\frac{1}{2}\lVert\omega^{n+1}\rVert_{H^{2}}^{2}=\langle\partial_{t}J^{2}\omega^{n+1},J^{2}\omega^{n+1}\rangle_{2}$
$\langle J^{2}(u^{n}\cdot\nabla\omega^{n+1})-u^{n}\cdot\nabla J^{2}\omega^{n+1},J^{2}\omega^{n+1}\rangle_{2}-\langle u^{n}\cdot\nabla J^{2}\omega^{n+1},J^{2}\omega^{n+1}\rangle_{2}+\langle J^{2}(\{\frac{1}{2}\lvert u^{n}\rvert^{2},\rho^{n}\}),J^{2}\omega^{n+1}\rangle_{2}$
Apply Kato-Ponce commutator estimate (cf. [KP88]) on the first term, integrating the second term by part and deleting it using divergence-free property of $u^{n}$, and using the algebra property of $H^{2}$ we conclude
$\frac{d}{dt}\frac{1}{2}\lVert\omega^{n+1}\rVert_{H^{2}}^{2}\lesssim\lVert\nabla u^{n}\rVert_{H^{2}}\lVert\omega^{n+1}\rVert_{H^{2}}^{2}+\lVert u^{n}\rVert_{H^{3}}\lVert\rho^{n}\rVert_{H^{2}}\lVert\omega^{n+1}\rVert_{H^{2}}$
which, together with our assumption $(\omega^{n},u^{n},\rho^{n})\in\mathcal{U}_{R}$, gives:
$\frac{d}{dt}\lVert\omega^{n+1}\rVert_{H^{2}}\lesssim R_{2}\lVert\omega^{n+1}\rVert_{H^{2}}+R_{2}R_{3}$
Gronwall inequality therefore implies:
$\lVert\omega^{n+1}(t)\rVert_{H^{2}}\leq\exp(CR_{2}t)(\lVert\omega_{0}\rVert_{H^{2}}+CR_{2}R_{3}t)$
@@
${CR_{2}T}\leq 1/2,\quad\frac{1}{2}\sqrt{e}R_{3}\leq\lVert\omega_{0}\rVert_{H^{2}}$
@@
$\lVert\rho^{n+1}\rVert_{H^{3}},\lVert 1/\rho^{n+1}\rVert_{H^{3}}\leq\sqrt{e}C(\rho_{0}),\quad\lVert\omega_{t}^{n+1}\rVert_{H^{2}}\leq 2\sqrt{e}\lVert\omega_{0}\rVert_{H^{2}}$
Hence we can choose $R_{1}=2\sqrt{e}\lVert\omega_{0}\rVert_{H^{2}}$.
Now we estimate $u^{n+1}$. Given $\lVert\omega^{n+1}\rVert\leq 2\sqrt{e}\lVert\omega_{0}\rVert_{H^{2}}$ and $\lVert\rho^{n+1}\rVert_{H^{3}}\leq\sqrt{e}C(\rho_{0})$, we have:
$\lVert u^{n+1}\rVert_{H^{3}}\lesssim\lVert\mathbf{L}_{\rho^{n+1}}^{-1}\omega^{n+1}\rVert_{H^{4}}=\lVert\psi^{n+1}\rVert_{H^{4}}\leq 2\sqrt{e}CP(\sqrt{e}C(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}}$
@@
$R_{2}=2\sqrt{e}CP(\sqrt{e}C(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}},\quad R_{3}=\sqrt{e}C(\rho_{0})$
@@
Denote by $C_{R}$ constants depending on $R_{1},R_{2},R_{3}$. We now again make use of the (51) to control the RHS:
@@
$\lVert\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}}$
@@
$\lVert\delta u^{n}\rVert_{H^{2}}\leq\lVert\psi^{n+1}-\psi^{n}\rVert_{H^{3}}\leq C_{R}(\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}})$ (54)
@@
Again, applying $J$, $J^{2}$ to above two equations, pairing with $J\delta\omega^{n}$, $J^{2}\delta\rho^{n}$ respectively, thanks to div-free property of $u^{n}$ we have
$\frac{d}{dt}\frac{1}{2}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+\langle J(u^{n}\cdot\nabla\delta\omega^{n})-u^{n}\cdot\nabla J\delta\omega^{n},J\delta\omega^{n}\rangle_{2}+\langle J\nabla\cdot(\delta u^{n}\omega^{n}),J\delta\omega^{n}\rangle_{2}=\langle J\sigma^{n},J\delta\omega^{n}\rangle$
$\frac{d}{dt}\frac{1}{2}\lVert\delta\rho^{n}\rVert_{H^{2}}^{2}+\langle J^{2}(u^{n}\cdot\nabla\delta\rho^{n})-u^{n}\cdot\nabla J^{2}\delta\rho^{n},J^{2}\delta\rho^{n}\rangle_{2}+\langle J^{2}(\delta u^{n}\cdot\nabla\rho^{n}),J^{2}\delta\rho^{n}\rangle_{2}=0$
Now, thanks to Kato-Ponce and the algebra property of $H^{2}$, $H^{3}$, we conclude
$\frac{d}{dt}\frac{1}{2}\lVert\delta\rho^{n}\rVert_{H^{2}}^{2}\leq C_{R}\lVert\delta\rho^{n}\rVert_{H^{2}}^{2}+C_{R}\lVert\delta u^{n}\rVert_{H^{2}}\lVert\delta\rho^{n}\rVert_{H^{2}}$ (55)
$\frac{d}{dt}\frac{1}{2}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}\leq C_{R}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+C_{R}\lVert\delta u^{n}\rVert_{H^{2}}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\sigma^{n}\rVert_{H^{1}}\lVert\delta\omega^{n}\rVert_{H^{1}}$
Notice that
@@
$\frac{d}{dt}\frac{1}{2}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}\leq C_{R}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+C_{R}\lVert\delta u^{n}\rVert_{H^{2}}\lVert\delta\omega^{n}\rVert_{H^{1}}+C_{R}(\lVert\delta\rho^{n-1}\rVert_{H^{2}}+\lVert\delta u^{n-1}\rVert_{H^{2}})\lVert\delta\omega^{n}\rVert_{H^{1}}$ (56)
Step 3. Now we define: $E^{n}(t):=\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+\lVert\delta\rho^{n}\rVert_{H^{1}}^{2}$. By (54), (55), (56), we obtain:
$\frac{d}{dt}E^{n}(t)\leq C_{R}E^{n}(t)+C_{R}E^{n-1}(t),\quad E(0)=0$
Now, integrating above inequality from time 0 to $t$, and denote by $U_{n}:=\sup_{t\in[0,T]}E^{n}(t)$. We have:
@@
$U_{n}=\sup_{t\in[0,T]}\lVert\delta\omega^{n}\rVert_{H^{1}}^{2}+\lVert\delta\rho^{n}\rVert_{H^{1}}^{2}.$

### v2

@@ Acknowledgements
The author would like to thank V.Vicol for stimulating discussions which initiate the study of Lagrangian analyticity of IIE. The author thanks T.Drivas for pointing out the ongoing work [DG26+] and for helpful discussions. The author thanks B.Khesin, A.Mazzucato and A.Novikov for stimulating discussions and valuable suggestions. This work was partially supported by the US National Science Foundation grants DMS-1909103 and DMS-2206453 (PI Mazzucato).
@@ 2 Geometry of $\mathsf{SDiff}(D)$ with Weighted Metric
In this section, we study the Riemannian structure of $\mathsf{SDiff}(D)$ which naturally fits the inhomogeneous incompressible Euler equation. Our presentation in this section mainly follows and generalizes the discussions in the unpublished notes [DrWS25], which generalizes constrained geodesic, second fundamental form and sectional curvature of level set submanifold from finite to infinite dimension (The $\mathsf{SDiff}(D)$ setting).
We start with the well-known Jacobi-Liouville formula which will be repeatedly used in this section.
@@
Remark 2.2
As pointed out by [Kh25], yet another simpler proof of the above lemma is by Abel’s theorem on Wronskian of matrix ODEs. More precisely, $\det A$ is the Wronskian of matrix $A$, which satisfies $\dot{A}=AA^{-1}\dot{A}$, then apply Abel’s theorem to coefficient matrix $\dot{A}A^{-1}$ the desired result follows.
@@ 2.1 The Riemannian Structure
Let $D=\mathbb{T}^{d}$, which is the physical domain of our fluids, equipped with natural periodic boundary conditions. Throughout this paper, we denote by $\mathsf{Diff}(D)$($\mathsf{SDiff}(D)$) the Lie group of (volume-preserving) diffeomorphisms in $D$. We denote by $\mathfrak{X}(D)$ ($\mathfrak{X}_{\sigma}(D)$) the space of (divergence-free) smooth vector fields on $D$. We remark that the formulation below also works in the setting of smooth compact Riemannian manifold, and works under $H^{s+1}$ Sobolev regularity with $s>d/2$, after suitable technical modification. Here, we only consider the smooth case on torus for simplicity.
@@
here $\rho=X_{\sharp}\rho_{0}=\rho_{0}\circ A$. Clearly $\rho$ is a rearrangement of $\rho_{0}$, so the assumption (A) is preserved under the action of $\mathsf{SDiff}(D)$ on $\mathsf{Dens}(D)$.
@@
We now extend the domain of the above normal generator $\mathcal{N}$ to distributions. Let $\mathcal{M}$ be the set of Borel measures on $D$ with finite variation, for any $\mu\in\mathcal{M}$ we define the following co-normal functional
@@
which satisfies the annihilating property $T_{X}\mathsf{SDiff}(D)\subset\mathsf{Ker}(\mathcal{N}_{X}^{*}(\mu))$.
@@
    where $\Delta_{\rho}=\nabla\cdot(\rho^{-1}\nabla)$ and $G_{\rho}$ is the Green’s function of the elliptic operator $\Delta_{\rho}$, under mean-zero condition.
@@ 2.3 Second Fundamental Form and Curvature Tensor
$\int_{D}[\mathbf{Q}_{\rho}(u\cdot\nabla u)\mathbf{Q}_{\rho}(v\cdot\nabla v)-\lvert\mathbf{Q}_{\rho}(u\cdot\nabla v)\rvert^{2}]\rho(z)dz.$
@@
$\ddot{X}(t,a)=(\partial_{t}u+u\cdot\nabla u)(X_{t}(a))=0.$
which implies that particle trajectories are all straight lines.
@@ 2.4 A Gibbs-Appell Formulation
Different from the celebrated Lagrange’s least action principle and Hamilton’s stationary action principle in analytical mechanics, there is a much lesser popular principle/formalism in classic mechanics, known as Gibbs-Appell formulation, which connects to the so-called Gauss’s principle of least curvature and Hertz’s principle of least constraint (see [[Lew96]]). The natural playground for this formulation is a mechanical system with ideal constraints, which is compatible with the nature of incompressible fluids. We refer to [[TG23]] for the study of the Navier-Stokes equation from this point of view. Instead of minimizing Lagrangian action under the constraint of fixed endpoints, the Gibbs-Appell principle minimizes the so-called Appellian functional, which is a quadratic function in acceleration, at each instant.
@@
for $X\in C^{2}([0,T];\mathsf{Diff}(D))$. Then minimizer $X_{*}$ of $\mathscr{A}$ at each instant $t\in[0,T]$, subject to the constraint $X(t)\in\mathsf{SDiff}(D)$, is the solution of the following Newton’s equation:
@@ 3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
Let $D=\mathbb{R}^{d}$ or $\mathbb{T}^{d}$, we now state the following:
@@
    Given initial data $(u_{0},\xi_{0})$ satisfying compatibility condition, the solution of (30) is given by the following Lagrangian-Eulerian system:
@@ 3.2 Vorticity Formulation
We formally unify the above relations in $2$-d in the following commutative diagram.
@@
$\mathbf{P}_{\rho}(\rho^{-1}\eta)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times\eta.$
@@
Assume $(u,\rho)$ is a smooth solution of IIE, then for any regular value $\alpha>0$ of $\rho_{0}$, where $E_{\alpha}:=\{\rho_{0}=\alpha\}$ is a smooth loop, let $D_{\alpha}:=\{\rho_{0}<\alpha\}$ be the smooth bounded domain enclosed by $E_{\alpha}$, then we have
@@ 3.3 Local Existence Via Vorticity Formulation
Assume $\omega\in H_{0}^{2}$, $\rho\in H^{3}$ satisfying the assumption (A), then any mean-zero weak solution $\psi\in\dot{H}_{\rho}^{1}$ of $\mathbf{L}_{\rho}\psi=\omega$ also satisfies $\psi\in H_{0}^{4}$.
@@
Step 1.($H^{1}$ Estimate) Multiply both sides of the equation and integrating by part, we have:
@@
Now, since $\rho\in H^{3}$ and the function $f(z)=z^{-1}$ is smooth on the range of $\rho$, we have $\rho^{-1}=f(\rho)\in H^{3}$. More precisely, we have
$\lVert\rho^{-1}\rVert_{H^{2}}=\sum_{j=0}^{2}\lVert D^{j}(1/\rho)\rVert_{2}\lesssim\sum_{j=0}^{2}\lVert\rho\rVert_{H^{2}}^{j}$
Hence, we conclude:
$\lVert\psi\rVert_{H^{3}}\lesssim\lVert\omega\rVert_{H^{1}}\lVert\rho^{-1}\rVert_{H^{2}}(1+\lVert\rho\rVert_{H^{3}}+\lVert\rho\rVert_{H^{3}}^{2})\lesssim\lVert\omega\rVert_{H^{1}}\sum_{j=1}^{4}\lVert\rho\rVert_{H^{3}}^{j}$
@@
where $P$ is some polynomial with positive coefficients. Hence, it is monotonically increasing on $(0,\infty)$. ∎
We now prove the local existence of IIE using our vorticity formulation, for simplicity we consider the casr $d=2$. Similar idea can be applied directly to the system (1) to conclude local well-posedness of strong solution in all dimensions. For a similar Lagrangian approach to the local well-posedness of hydrodynamic models, we refer the interested readers to [CI08][MP25][Zh10].
@@
Assume $d=2$, Let $(\omega_{0},\rho_{0})\in H^{2}\times H^{3}$ and let $\rho$ satisfies Assumption (A), then (47) admits a unique solution local in time.
@@
Now we consider the estimate for vorticity equation. By Duhamel’s principle, we can write down the following Lagrangian solution formulae:
$\omega^{n+1}(t,\cdot)=\omega_{0}\circ A_{t}^{n}-\bigg[\int_{0}^{t}(\nabla^{*}u^{n}u^{n}\cdot\nabla^{\perp}\rho^{n})(X_{\tau}^{n})d\tau\bigg]\circ A_{t}^{n}$
$\omega_{0}\circ A_{t}^{n}-\bigg[\nabla^{\perp}\rho_{0}\int_{0}^{t}\nabla^{*}X^{n}(\nabla^{*}u^{n}u^{n})(X_{\tau}^{n})d\tau\bigg]\circ A_{t}^{n}$
Hence, we may appeal to the algebra structure of $H^{2}$ and tandard result in propagation of Sobolev regularity along smooth flow (for instance, see [BCD11][MP25]) to estimate:
$\lVert\omega^{n+1}\rVert_{H^{2}}\leq\lVert\omega_{0}\circ A_{t}\rVert_{H^{2}}+\lVert\nabla\rho_{0}\circ A_{t}\rVert_{H^{2}}\exp(CtR_{2})\int_{0}^{t}\lVert\nabla X^{n}\rVert_{H^{2}}\lVert(\nabla^{*}uu)(X^{n})\rVert_{H^{2}}d\tau$
$\leq\lVert\omega_{0}\rVert_{H^{2}}\exp(CtR_{2})+R_{3}\exp(2CtR_{2})\int_{0}^{t}\exp(2C\tau R_{2})R_{2}^{2}dt$
$\leq\exp(CtR_{2})(\lVert\omega_{0}\rVert_{H^{2}}+CR_{2}^{2}R_{3}t)$ (54)
@@
${CR_{2}T}\leq 1,\quad CR_{2}^{2}R_{3}T\leq\lVert\omega_{0}\rVert_{H^{2}}$
@@
$\lVert\rho^{n+1}\rVert_{H^{3}},\lVert 1/\rho^{n+1}\rVert_{H^{3}}\leq eC(\rho_{0}),\quad\lVert\omega_{t}^{n+1}\rVert_{H^{2}}\leq 2e\lVert\omega_{0}\rVert_{H^{2}}$
Hence, we can choose $R_{1}=2e\lVert\omega_{0}\rVert_{H^{2}}$.
Now we estimate $u^{n+1}$. Given $\lVert\omega^{n+1}\rVert\leq 2e\lVert\omega_{0}\rVert_{H^{2}}$ and $\lVert\rho^{n+1}\rVert_{H^{3}}\leq eC(\rho_{0})$, we have:
$\lVert u^{n+1}\rVert_{H^{3}}\lesssim\lVert\mathbf{L}_{\rho^{n+1}}^{-1}\omega^{n+1}\rVert_{H^{4}}=\lVert\psi^{n+1}\rVert_{H^{4}}\leq 2eCP(eC(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}}$
@@
$R_{2}=2eCP(eC(\rho_{0}))\lVert\omega_{0}\rVert_{H^{2}},\quad R_{3}=eC(\rho_{0}).$
@@
Denote by $C_{R}$ constants depending on $R_{1},R_{2},R_{3}$ which may change from line to line. We now again make use of the (51) to control the RHS:
@@
$\lVert\nabla(\log\rho^{n+1}-\log\rho^{n})\nabla\psi^{n}\rVert_{H^{1}}\leq C_{R}\lVert\rho^{n+1}-\rho^{n}\rVert_{H^{2}}.$
@@
$\lVert\delta u^{n}\rVert_{H^{2}}\leq\lVert\psi^{n+1}-\psi^{n}\rVert_{H^{3}}\leq C_{R}(\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}).$ (55)
@@
Again, using Duhamel’s formula we conclude:
$\delta\omega^{n}=\delta\omega_{0}\circ A_{t}^{n}+\bigg(\int_{0}^{t}(\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n}$
$\delta\rho^{n}=\delta\rho_{0}\circ A_{t}^{n}-\bigg(\int_{0}^{t}(\delta u^{n}\cdot\nabla\rho^{n})(X_{\tau}^{n})d\tau\bigg)\circ A_{t}^{n}$
Follow the same line as for $\omega^{n}$, since $\delta\omega_{0}=\delta\rho_{0}=0$, we have the following Sobolev estimates:
$\lVert\delta\omega^{n}\rVert_{H^{1}}\leq e\int_{0}^{t}\lVert\sigma_{n}-\delta u^{n}\cdot\nabla\omega^{n}\rVert_{H^{1}}d\tau\leq e\int_{0}^{t}\lVert\sigma_{n}\rVert_{H^{1}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau$
$\lVert\delta\rho^{n}\rVert_{H^{2}}\leq e\int_{0}^{t}\lVert\delta u^{n}\cdot\nabla\rho^{n}\rVert_{H^{2}}d\tau\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau$ (56)
Now we estimate $\sigma_{n}$. Notice that
@@
$\lVert\sigma_{n}\rVert_{H^{1}}\leq C_{R}(\lVert\delta\omega^{n-1}\rVert_{H^{1}}+\lVert\delta\rho^{n-1}\rVert_{H^{2}})$
and
$\lVert\delta\omega^{n}\rVert_{H^{1}}\leq C_{R}\int_{0}^{t}\lVert\delta\omega^{n-1}\rVert_{H^{1}}+\lVert\delta\rho^{n-1}\rVert_{H^{2}}d\tau+C_{R}\int_{0}^{t}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{2}}d\tau$ (57)
Step 3. Now we define: $E^{n}(t):=\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{1}}$. By (55), (56), (57), we obtain:
$E^{n}(t)\leq C_{R}\int_{0}^{t}E^{n}(\tau)d\tau+C_{R}\int_{0}^{t}E^{n-1}(\tau)d\tau,\quad E(0)=0$
Now we denote by $U_{n}:=\sup_{t\in[0,T]}E^{n}(t)$ and rewrite the above inequality as
@@
$U_{n}=\sup_{t\in[0,T]}\lVert\delta\omega^{n}\rVert_{H^{1}}+\lVert\delta\rho^{n}\rVert_{H^{1}}.$